# Assessment of Spatio-Temporal Predictors in the Presence of Missing and Heterogeneous Data

### v2

@@ 6.3 Analysis with missing and heterogeneous data
Lastly, we report the scores of the residual analysis carried out on similar synthetic data where, however, about 20% of the observations were missing and where the data comes from heterogeneous sensors. For the scores of these experiments, we refer the reader to Figures 7 and 8.
@@
Despite the presence of data gaps, visually represented as white areas within the representation of $c_{\lambda=1/2}(t,v)$ in Figure 7, the reported correlation scores exhibit patterns aligned with those present in the data and discussed in previous Section 6.2.
To validate the effectiveness of the proposed analysis on data coming from heterogeneous sensors, the residuals $\boldsymbol{\mathbf{r}}$ are generated following the process described in Section 6.1, but starting from non-i.i.d. noise $\varepsilon$. In particular, each noise component $\varepsilon_{t,v}$ is sampled from either of three zero-median distributions: a uniform distribution, a Laplace distribution, a bimodal distribution generated from a mixture of 2 Gaussian distributions; the distribution assignment is uniform but random.
@@
The analysis reported in Figure 8 shows scores very similar to those of Figures 5 and 7. In particular, both regions $A$ and $B$ are clearly identifiable and analog conclusions can be drawn.
To conclude, the reported results validate the proposed AZ-analysis by showing that both spatial and temporal correlation can be identified. In particular, it can be identified at a global level (Q1), at the level of single nodes (Q2) and of time steps (Q3). Secondly, the results confirm the broad applicability of the presented residual analysis which can effectively operate under mild assumptions, and in presence of missing and heterogeneous data.
@@ 7.1 Data and models
The selected set of predictive models includes generic baselines for time series forecasting and well-established methods from the literature: GWNet (Wu et al., 2019), DCRNN (Li et al., 2018), and AGCRN (Bai et al., 2020). The baseline models consist of a recurrent neural network (RNN) with gated recurrent units (GRUs) trained and operated on individual univariate time series (uvRNN), a second RNN trained and operated on the multivariate time series obtained by stacking all univariate time series (mvRNN), and a generic time-then-space STGNN (ttsRNN) processing the output of uvRNN with a graph neural network (Cini et al., 2023b). Additional experimental details are provided in Appendices B and C.
@@
Score $c_{\lambda=1/2}(t)$ Refer to caption
MAE Refer to caption
@@
Score $c_{\lambda=1/2}(t)$ Refer to caption
Figure 9: Left) Temporal scores $c_{\lambda=1/2}(t)$ of the residuals (top) and predictive MAE (bottom) on MetrLA. Shaded areas highlight regions of imputed target data. Right) Temporal scores $c_{\lambda=1/2}(t)$ on 1-step ahead prediction on MetrLA. Shaded areas highlight regions of imputed target data.
@@ 7.2 Collected insights and discussion
Figure 9(a) presents the time scores $c_{\lambda=1/2}(t)$ and prediction error (expressed here in terms of MAE) for all considered models. For visualization purposes, a window of 500 time steps and 40 nodes is depicted in the figures.
@@
From the top part of Figure 9(a), we observe that regardless of the predictive model, the time scores are all consistently high within two timeframes around $t=100$ and $t=450$.
As indicated by the shaded areas in the figure, these regions correspond to time steps with artificially imputed data. The observed high scores likely relate with the imputation method employed, which replicates the last observed value.
@@
Comparing Figure 9(a), which computes time scores from multi-step ahead predictions, with Figure 9(b) focusing on 1-step ahead predictions, reveals that the correlation patterns are more pronounced in the latter case. Furthermore, the observed correlation is diminished in regions where data is not imputed, suggesting that there is less room for improvement in short-term predictions compared to long-term predictions.
Extending the residual analysis for one of the models, GWNet, yields further relevant insights; please, refer to Figure 10.
@@
Scores $c_{\lambda}(t,v)$ confirm that the aforementioned timeframes are indeed problematic. Moreover, they highlight specific nodes, namely nodes 28 and 29, that warrant closer attention. Further examination of the mask of imputed data at the bottom of Figure 10 reveals that the time series for these nodes have also been imputed.
@@ 8.1 Data and models
We consider EngRAD dataset (Marisca et al., 2024), which encompasses 5 meteorological variables related to photovoltaic energy production. Hourly estimations of these variables are obtained from a grid of 487 locations across England. A graph representation is constructed based on the pairwise geographical distance between locations. The ttsRNN model was employed and trained to minimize the MAE on 3-hour ahead forecasts of horizontal irradiance, leveraging an input window of 24 preceding hours and related weather variables. The residual analysis was conducted on the test set, focusing on 1-hour ahead forecasts. Further experimental details are provided in Appendices B and C.
@@ 8.2 Insights from the residual analysis
Figure 11 reveals a daily trend evident in both the MAE and the time scores $c_{\lambda}(t)$, attributable to the natural solar irradiation cycle. Missing observations correspond to the negligible solar irradiation of the night.
@@
Higher correlations observed during dawn and dusk suggest room for model improvement, while the MAE remains relatively low throughout these transitional periods. While the MAE alone is not particularly concerning during dawn and dusk, further inspection reveals higher relative prediction error (mean absolute percentage error, MAPE), reinforcing the evidence provided by $c_{\lambda}(t)$ that the model’s predictions can improve in these hours of the day.
Furthermore, the residual analysis unveils another interesting insight: in certain instances where the model exhibits low MAE values, a high correlation is present, thus indicating that they can be further enhanced. Specifically, on June 23rd and 24th, 2020, the target variable exhibited a notably high magnitude with reduced variability, coupled with a consistently lower MAE. While this suggests that the target variable is easier to forecast during these specific periods, it is likely that the training primarily focused on regions characterized by higher prediction errors.
@@ 9 Conclusions
In this paper, we introduce a novel analysis of model prediction residuals, called AZ-analysis, for spatio-temporal data that allows practitioners to identify space-time regions, such as time intervals or sets of sensors, where the model performance can be improved. The proposed analysis complements that of the prediction error by evaluating the presence of temporal and spatial residual correlation in regions of the space-time and sheds light on aspects where the predictive model is not optimal. The AZ-analysis is well-suited for assessing deep neural models in real-world scenarios as (i) it can operate in presence of missing heterogeneous data, (ii) it scales to large and complex datasets by exploiting graph-side information, and (iii) only requests residuals centered around zero in a few setups; no other assumption about the distribution of residuals is requested, not even of identical distribution. The proposed framework is first validated on synthetic experiments and general comments on how to interpret the results are provided along with real-world use cases in the context of forecasting traffic flow and energy production to showcase the potential of the proposed AZ-analysis.
@@ Appendix A Experiment on the comparability of scores
Dimension $d$ of residual vectors is set to 5. Mean, standard deviation, and 25% and 75% percentiles are reported in Figure 3 and are estimated over 100 repeated simulations.
@@ Appendix B Additional details on models and training
For the traffic forecasting task, predictions are made from an input window of 12 time steps (1 hour), with the hour of the day and the day of the week added as covariates to all models. For the solar irradiance prediction, forecasts are issued for 6 hours ahead with an input window of 24 hours; here, the hour of the day and hour of the year are added as covariates. The residual analyses are conducted on the test set considering traffic forecasts at 5, 20, 40 and 60 minutes and solar irradiance predictions in the next hour for the energy production.
Baseline models uvRNN, mvRNN and ttsRNN are 1-layer RNN with gated recurrent units and the exponential linear unit as activation function. Each model has 32 hidden neurons. ttsRNN implements a simple message passing based on average aggregation on the uvRNN outputs. Models are trained to minimize the mean absolute error (MAE) on the first 70% of data, validated on the following 10%, and tested on the remaining 20%; missing and imputed data are appropriately masked out during training. Training is performed for 200 epochs using Adam optimizer with an initial learning rate of 0.003 decreased every 50 epochs by a factor of 4. Each epoch consists of 300 batches of size 64. Early stopping is triggered after 50 epochs of no improvement on the validation set.
Appendix C Hardware and software
Model training and inference are performed on a workstation equipped with Intel(R) Xeon(R) Silver 4116 CPU 2.10GHz processors and NVIDIA TITAN V GPUs. The analysis of residuals does not require GPU acceleration. Experiments are developed in Python 3.8 mainly relying on open-source libraries PyTorch (Paszke et al., 2019), PyTorch Geometric (Fey and Lenssen, 2019), Torch Spatiotemporal (Cini and Marisca, 2022) and NumPy (Harris et al., 2020). The source code to reproduce the results will be released upon publication.

### v3

@@ 6.3 Analysis with missing and heterogeneous data
Lastly, we report the scores of the residual analysis carried out on similar synthetic data in which about 20% of the observations are missing and the data come from heterogeneous sensors. For the scores of these experiments, we refer the reader to Figures 8 and 9.
@@
Despite the presence of data gaps, visually represented as white areas within the representation of $c_{\lambda=1/2}(t,v)$ in Figure 8, the reported correlation scores exhibit patterns aligned with those present in the data and discussed in the previous Section 6.2.
To validate the effectiveness of the proposed analysis on data coming from heterogeneous sensors, the residuals $\boldsymbol{\mathbf{r}}$ are generated following the process described in Section 6.1, but starting from non-i.i.d. noise $\varepsilon$. In particular, each noise component $\varepsilon_{t,v}$ is sampled from one of three zero-median distributions: a uniform distribution, a Laplace distribution, a bimodal distribution generated from a mixture of two Gaussian distributions; the distribution assignment is uniform at random.
@@
The analysis reported in Figure 9 shows scores very similar to those of Figures 6 and 8. In particular, both regions $A$ and $B$ are clearly identifiable and analogous conclusions can be drawn.
We now examine how our scores compare to alternative correlation measures commonly used in temporal and spatial analysis, which provide natural points of reference for assessing the behavior of the proposed scores: $k$-lag autocorrelation and Moran’s $I$ statistics (Moran, 1950; Anselin, 1995). For each node $v$, we consider the aggregated autocorrelation measure
$\text{Autocorr}_{v}(\bar{k})\doteq\frac{1}{\bar{k}}\sum_{k=1}^{\bar{k}}\rho_{k}(v),$ (32)
where $\rho_{k}(v)$ estimates the autocorrelation $\text{corr}(r_{v,t},r_{v,t-k})$ at lag $k>0$. This provides a summary of temporal dependence up to lag $\bar{k}$. For spatial dependence, we consider Moran’s $I$ statistic computed at each time step $t$
$I_{t}\doteq\frac{\sum_{(u,v)\in E_{t}}w_{t,(u,v)}\,(r_{t,u}-\bar{r}_{t})\,(r_{t,v}-\bar{r}_{t})}{\lVert\boldsymbol{\mathbf{w}}_{t}\rVert_{1}\;s_{t}^{2}},$ (33)
and the Local Moran’s $I$ statistic to capture local spatial patterns
$I_{t,v}\doteq\frac{\sum_{u\in N_{t}(v)}w_{t,(u,v)}\,(r_{t,u}-\bar{r}_{t})\,(r_{t,v}-\bar{r}_{t})}{\rVert\{w_{t,(u,v)}:u\in N_{t}(v)\}\rVert_{1}\;s_{t}^{2}},$ (34)
where $\bar{r}_{t}$ denotes the mean residual at time $t$ and $s_{t}^{2}=|V_{t}|^{-1}\sum_{v\in V_{t}}(r_{t,v}-\bar{r}_{t})^{2}$.
Comparison with other correlation analyses
Figure 10 shows that $c_{0}(v)$ behaves similarly to Autocorr($k$): both increase in regions with temporal correlation. Autocorr(10) takes lower values than Autocorr(1) because the data exhibit only 1-lag dependence. For spatial dependence, the time scores $c_{1}(t)$ closely match the behavior of Moran’s $I$ and identify the same interval of spatial correlation. At the local level, both the proposed local scores $c_{\lambda}(t,v)$ and the Local Moran’s $I$ detect region $A$ as spatially correlated, with $c_{\lambda}(t,v)$ computed over 4-hop neighborhoods doing so more distinctly.
To conclude, the reported results validate the proposed AZ-analysis by showing that both spatial and temporal correlations can be identified. In particular, they can be identified at a global level (Q1), at the level of single nodes (Q2), and at the level of time steps (Q3). Moreover, the AZ-analysis identifies patterns aligned with existing approaches designed for either spatial or temporal correlation, while enabling a comprehensive spatio-temporal analysis of residual correlation under mild assumptions. The results also confirm the broad applicability of the proposed residual analysis, which can effectively operate under mild assumptions and with missing and heterogeneous data.
@@ 7.1 Data and models
The selected set of predictive models includes generic baselines for time series forecasting and well-established methods from the literature: GWNet (Wu et al., 2019), DCRNN (Li et al., 2018), and AGCRN (Bai et al., 2020). The baseline models consist of a recurrent neural network (RNN) with gated recurrent units (GRUs) trained and operated on individual univariate time series (uvRNN), a second RNN trained and operated on the multivariate time series obtained by stacking all univariate time series (mvRNN), and a generic time-then-space STGNN (ttsRNN) processing the output of uvRNN with a graph neural network (Cini et al., 2023a). Additional experimental details are provided in Appendices B and C.
@@
Figure 11: Left) Time scores $c_{\lambda=1/2}(t)$ of the residuals (top) and predictive MAE (bottom) on MetrLA. Right) Time scores $c_{\lambda=1/2}(t)$ on 1-step ahead prediction on MetrLA. The shaded area in gray represents the min-max range of scores/MAE obtained from 18 different models (3 runs for each listed model). Time frames highlighted in yellow denote imputed target data.
Figure 12: Node scores $c_{\lambda=0}(v)$ of the residuals (left) and predictive MAE (right) on MetrLA. The gray shaded band shows the min–max range across 18 model instances (three runs for each listed model). Nodes highlighted in blue are those where temporal correlation was artificially injected at test time.
@@ 7.2 Collected insights and discussion
Figure 11 presents the time scores $c_{\lambda=1/2}(t)$ and prediction error (expressed here in terms of MAE) for all considered models. For visualization purposes, a window of 500 time steps and 40 nodes is depicted in the figures.
@@
From the top part of Figure 11(a), we observe that regardless of the predictive model, the time scores are all consistently high within two time intervals around $t=100$ and $t=450$; as indicated by the shaded areas in the figure, these regions correspond to time steps with artificially imputed data. The observed high scores likely relate to the imputation method employed, which replicates the last observed value.
@@
Comparing Figure 11(a), which computes time scores from multi-step ahead predictions, with Figure 11(b) focusing on 1-step ahead predictions, reveals that the correlation patterns are more pronounced in the latter case. Furthermore, the observed correlation is diminished outside the imputed regions, suggesting that there is less room for improvement in short-term predictions compared to long-term predictions.
To corroborate the ability of the proposed scores to identify correlation, we manually introduced temporal dependence in the test data for five nodes highlighted in blue in Figure 12. Specifically, we enforced correlation by applying a moving average of width three: $\widetilde{\boldsymbol{\mathbf{x}}}_{t,v}=(\boldsymbol{\mathbf{x}}_{t-1,v}+\boldsymbol{\mathbf{x}}_{t,v}+\boldsymbol{\mathbf{x}}_{t+1,v})/3$. Figure 12 reports the resulting node scores and prediction MAE across all considered models. The injected correlation is clearly reflected by the proposed node scores $c_{\lambda=0}(v)$, whereas the MAE appears essentially unaffected.
Extending the residual analysis for one of the models, GWNet, yields further relevant insights; refer to Figure 13.
@@
Scores $c_{\lambda}(t,v)$ confirm that the aforementioned time frames are indeed problematic. Moreover, they highlight specific nodes, namely nodes 28 and 29, that warrant closer inspection. Further examination of the mask of imputed data at the bottom of Figure 13 reveals that the time series for these nodes have also been imputed.
@@
(a) uvRNN
(b) DCRNN
(c) AGCRN
(d) GWNet
Figure 15: Analysis of the prediction residuals for energy production prediction, extending Figure 14 to other models.
@@ 8.1 Data and models
We consider the EngRAD dataset (Marisca et al., 2024), which encompasses 5 meteorological variables related to photovoltaic energy production. Hourly estimates of these variables are obtained from a grid of 487 locations across England. A graph representation is constructed based on the pairwise geographical distance between locations. The ttsRNN model was employed and trained to minimize the MAE on 3-hour-ahead forecasts of horizontal irradiance, leveraging an input window of 24 preceding hours and related weather variables. The residual analysis was conducted on the test set, focusing on 1-hour-ahead forecasts. Further experimental details are provided in B and C.
@@ 8.2 Insights from the residual analysis
Figure 14 reveals a clear daily trend in both the MAE and the time scores $c_{\lambda}(t)$, reflecting the natural solar irradiation cycle. Missing observations correspond to nighttime periods with negligible irradiance.
@@
Higher correlations observed during dawn and dusk suggest room for model improvement, while the MAE remains relatively low throughout these transitional periods. While the MAE alone is not particularly concerning during dawn and dusk, further inspection reveals a higher relative prediction error (mean absolute percentage error, MAPE), reinforcing the evidence provided by $c_{\lambda}(t)$ that the model’s predictions can improve in these hours of the day.
Beyond these effects, the residual analysis unveils an additional pattern: in certain instances where the model exhibits low MAE values, a high correlation is present, suggesting that the predictions could be further refined. For example, on June 23rd and 24th, 2020, the target variable attains a notably high magnitude with reduced variability, coupled with a consistently lower MAE. While this suggests that the target variable is easier to forecast during these specific periods, it is plausible that the model training is more strongly influenced by regions characterized by higher prediction errors. Analogous behavior is visible for different models too, as shown in Figure 15.
@@ 9 Conclusions
In this paper, we introduce a novel analysis of model prediction residuals, called AZ-analysis, for spatio-temporal data. The proposed approach allows practitioners to identify regions, such as time intervals or groups of sensors, where model performance can be improved. This analysis complements traditional prediction-error evaluations by assessing temporal and spatial correlation in the residuals. It helps reveal aspects of the predictive model that may be suboptimal. The AZ-analysis is well-suited for assessing deep neural models in real-world scenarios. It offers several advantages: (i) it can operate in the presence of missing and heterogeneous data; (ii) it scales to large and complex datasets by exploiting graph side information; and (iii) it requires only that residuals are centered around zero in a few specific setups. No further assumptions about the residuals’ distribution are necessary – not even identical distributions. We first validate the proposed framework through synthetic experiments and provide general guidance on how to interpret the results. We then showcase its potential through real-world use cases in forecasting traffic flow and energy production. Overall, AZ-analysis offers a powerful and widely applicable tool for gaining deeper insights into the behavior and limitations of spatio-temporal predictive models.
@@ Appendix A Experiment on the comparability of scores
Dimension $d$ of residual vectors is set to 5. Mean, standard deviation, and 25th and 75th percentiles are reported in Figure 3 and are estimated over 100 repeated simulations.
@@ Appendix B Additional details on models and training
For the traffic forecasting task, predictions are generated from an input window of 1 hour (12 time steps) to forecast the subsequent hour. The hour of the day and the day of the week are included as covariates for all models. For the solar irradiance task, forecasts are issued 6 hours ahead using an input window of 24 hours; in this case, the hour of the day and the hour of the year are included as covariates. Residual analyses are conducted on the test set considering traffic prediction horizons of 5, 20, 40, and 60 minutes, and 1-hour-ahead solar irradiance forecasts.
For the traffic forecasting task, predictions are made from an input window of 1 hour (12 time steps) to predict the subsequent hour. The hour of the day and the day of the week are included as covariates for all models. For the solar irradiance prediction, forecasts are issued 6 hours ahead using an input window of 24 hours; here, the hour of the day and hour of the year are added as covariates. The residual analyses are conducted on the test set considering traffic forecasts at 5, 20, 40, and 60 minutes, and 1-hour-ahead solar irradiance forecasts. The graphs are derived from pairwise geographical distances between sensors, which are processed through a Gaussian kernel to define edge weights. Edges with a weight below 0.1 are discarded; for EngRAD, connections beyond the 8 nearest neighbors are also removed.
Baseline models uvRNN, mvRNN, and ttsRNN are single-layer recurrent neural networks with gated recurrent units and exponential linear unit activation. uvRNN and mvRNN use 32 hidden units. ttsRNN uses 16 hidden units and implements a simple 2-layer message-passing scheme with average aggregation on the uvRNN outputs. All models are trained to minimize the mean absolute error (MAE), with missing and imputed data appropriately masked during training. For MetrLA, the first 70% of the data is used for training, the next 10% for validation, and the remaining 20% for testing. For EngRAD, the first two years are used for training, with 12 weeks across all four seasons reserved for validation, and the final year is used for testing. Training is performed for 200 epochs using the Adam optimizer with an initial learning rate of 0.003, reduced by a factor of 4 every 50 epochs. Each epoch consists of 300 batches of size 64. Early stopping is applied after 50 epochs without improvement on the validation set.
Appendix C Hardware, software and data
Model training and inference are performed on a workstation equipped with Intel(R) Xeon(R) Silver 4116 CPU 2.10GHz processors and NVIDIA TITAN V GPUs. The analysis of residuals does not require GPU acceleration. Experiments are developed in Python 3.8 mainly relying on open-source libraries PyTorch (Paszke et al., 2019), PyTorch Geometric (Fey and Lenssen, 2019), Torch Spatiotemporal (Cini and Marisca, 2022) and NumPy (Harris et al., 2020). In particular, Torch Spatiotemporal offers utilities for downloading the data used for the traffic and energy production use cases. Instructions for generating the synthetic residuals are detailed in the main paper. The source code to reproduce the results is available at: https://github.com/dzambon/az-analysis.